WHAT IS THE DIFFERENCE OF 8 1/2 AND 3 3/4 WRITTEN IN SIMPLEST TERMS?
Answer by checkley75(3666)

8 1/2=8.5
3 3/4=3.75 NOW FIND THE DIFFERENCE BY SUBTRACTING:
8.5-3.75=4.75 OR 4 3/4